OpenStudy (anonymous):

What are the vertex, focus, and directrix of the parabola with equation x=-3y^2? Please show all steps! Will give medal and fan for best answer!

OpenStudy (anonymous):

Do I plug -3y^2 into that..?

OpenStudy (anonymous):

uhm.. no

OpenStudy (anonymous):

what is (h,k) is this case?

OpenStudy (anonymous):

(x,y).. vertex.. I dont know. I'm confused with this. Sorry.

OpenStudy (anonymous):

http://www.purplemath.com/modules/parabola.htm

OpenStudy (anonymous):

x=-3y^2 <==> y^2 = (-1/3)x yes?

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

so what is the vertex in this case?

OpenStudy (anonymous):

0,0?

OpenStudy (anonymous):

correction: standard form should have been (y + k)^2 = 4 (p) (x - h)

OpenStudy (anonymous):

so how do we get 4 to show up in the equation (y - 0)^2 = (-1/3) (x - 0)?

OpenStudy (anonymous):

Put it in place of y?

OpenStudy (anonymous):

no, we can multiply the right side by (4/4) and that won't change the equation (y - 0)^2 = (4/4) (-1/3) (x - 0) (y - 0)^2 = (4) (1/4) (-1/3) (x - 0) (y - 0)^2 = (4) (-1/12) ((x - 0) so now what is the p?

OpenStudy (anonymous):

so the p is 4 then, right?

OpenStudy (anonymous):

or -1/12?

OpenStudy (anonymous):

yes p is -1/12. So what is the distance from the vertex to the focus?

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

no, it's 1/12

OpenStudy (anonymous):

That's what I meant lol.

OpenStudy (anonymous):

ok, so which way does the parabola opens?

OpenStudy (anonymous):

To the right..

OpenStudy (anonymous):

is p the directrix?

OpenStudy (anonymous):

no it opens to the left

OpenStudy (anonymous):

and no p is not the directrix, although we can use it to determine the directrix

OpenStudy (anonymous):

What does p stand for then? focus?

OpenStudy (anonymous):

p stands for the distance from the vertex to the focus

OpenStudy (anonymous):

which we found out to be 1/12

OpenStudy (anonymous):